OpenStudy (anonymous):

What does it mean for a number to be the "solution" to an equation? Give an example of a variable equation and its solution. Explain in complete sentences how you know that number is the solution.

OpenStudy (aaronq):

a number that is a solution to an equation is the value of a mathematical expression which corresponds to the parameters in question. eg. y=x^2 when x=2; y=2^2=4 using mathematical language, the expression y=x^2, means that y, the dependent variable, is a function of x, the independent variable, and so using arithmetic one can break down this function into y=x*x .. multiplication can be broken down into simplified counting. eg. when counting a grid.  one could count each individual row and add them together: 6+6+6+6+6+6+6+6 =48 multiplication is a simplified version of this: the number six eight times is forty eight. similarly, exponents are simplified expression for multiplications of the same number.. 6^2 .. 6 x 6 = 36